OpenStudy (anonymous):

Someone explain me, please f(x) = x sin x f'(x) = sin x + x cos x for some c, \(2\pi n 2\pi n cos(2\pi n+\dfrac {1}{n} \) why?

OpenStudy (anonymous):

@Supreme_Kurt

OpenStudy (welshfella):

the derivative of x sin x is found using the product rule (uv)' = uv' + vu' where u and v are functions of x

OpenStudy (welshfella):

the derivative of x is 1 and derivative of sin x is cos x

OpenStudy (anonymous):

@welshfella yes sir, :)

OpenStudy (triciaal):

not complete but the range of any angle, c, would be between 2 pi and 0

OpenStudy (anonymous):

I think he makes mistake. It should be \(2\pi~n<c<2\pi~n+\dfrac{1}{2}\) since we just focus on the left hand side, hence, ignore the right hand side

OpenStudy (anonymous):

\(2\pi n< c\\1<cosc\\2\pi~n<ccosc\\2\pi~n+ sin(2\pi n)<ccos c+sin c\\f'(c) > 2\pi n\)